OpenStudy (anonymous):

ken leans a 12-foot ladder against his house. He places the ladder so that the base is 5 feet from the house. How far up does the ladder reach?

OpenStudy (asadkarim7):

the pythagoras's theorm is being applied here hyp is 12 ft base is 5 ft sqrt 12^2- 5^2 = height 10.9 ft

OpenStudy (anonymous):

Use Pythgoras theorem here hypotenuse = 12 foot base = 5 ft perpendicular = ?

OpenStudy (anonymous):

a^2+b^2=c^2, the base is the 5 feet for b the ladder 12 is c, so then you figure out a